1. Use Pigeon hole principle to prove that any graph with at least 2 vertices contains two vertices of the same degree. (Hint: Prove by contradiction. (4 points)

2. Given 4n = (0)2 + (1)2 +...+)2- (6 Points)

a. Prove the above equation using binomial theorem. (3 Points)

b. Give a combinatorial proof for the given equating. (3 Points)
Answer #1

1.I assume we're talking about finite graphs. I'm pretty sure your statement is false for infinite graphs.

Assume that a finite graph G has n vertices. Then each vertex has a degree between n−1 and 0. But if any vertex has degree 0, then no vertex can have degree n−1, so it's not possible for the degrees of the graph's vertices to include both 0 and n−1. Thus, the n vertices of the graph can only have n−1 different degrees,

so by the pigeonhole principle at least two vertices must have the same degree.

2.(a)

By the binomial theorem, we know that we can write

(1+x)n=∑k=0n(nk)xk=(n0)+(n1)x+⋯+(nn)xn

we put x=1

2n=(n0)+(n1)+⋯+(nn)

4n=2n.2n

4n=2n.[(n0)+(n1)+⋯+(nn)]

4n=(n0)2n+(n1)2n+⋯+(nn)2n proved